consider the following short-run production function for a firm: show work

            Labor                           1          2          3          4          5          6          7          8

            Total product               5          12        ?          23        27        ?          32        33

            Average product          5          ?          6          5.75     5.4       5          ?          4.12

            Marginal product         5          7          6          ?          4          3          ?          1

The missing total product figures are

  1. 18 and 24
  2. 15 and 25
  3. 19 and 25
  4. 18 and 30
  5. None of the above
Answer #1

Answer

Option d

===========

total product of n input= total product of n-1 inputs+ marginal product of n th input

also, total product of n inputs =average product of n inputs * n

for Labor =3

total product is

from the second formula.

TP(3)=3*6=18 units

===

from the first formula

TP(3)=12+6=18 units

======================

TP(6)=5*6=30

TP(6)=27+3=30